Show that lim (|x| +1)= ∞
x→∞
1
Expert's answer
2021-08-04T16:41:20-0400

"\lim\limits_{x\to\infin}(|x|+1)=?\newline x\ is\ positive\ when\ x\to\infin.\ Therefore\ |x|=x.\newline =\lim\limits_x(x+1)\newline Applying\ Infinity\ property: \newline \lim\limits_{x\to\infin}(ax^{n}+...+bx+c)=\infin\newline a>0,\ n\ is\ odd\newline a=1,\ n=1\newline hence=\infin\newline thus\ \lim\limits_{x\to\infin}(|x|+1)=\infin"